OpenStudy (pounce0129):

Write an equation in point-slope form for the line that passes through the given point with the slope provided. (1,3) m= -3/4

OpenStudy (anonymous):

y-y1=m(x-x1) plug in your point (x1,y1) and the slope m

OpenStudy (pounce0129):

i did that and i changed it into slope intercept form, i got y=-3/4x + 3/2 and it just doesn't seem like the right answer

OpenStudy (anonymous):

Its asking for point slope form, not slope intercept form. When you plug in the values it will be in point slope form

OpenStudy (pounce0129):

my teacher asked us to change it into slope intercept form

OpenStudy (anonymous):

Oh, well then you would simplify it. y-3=-3/4(x-1) y-3=-3/4x+3/4 y=-3/4x+15/4

OpenStudy (pounce0129):

thanks